(Square root of 2X+3) = (square root of X+2) + 2
Can you help me solve this? I've already done some work:
1. I squared both sides of the equation and got: 2X+3 = X+2+(4 square root of X+2) + 4 because I foiled the right side, but I don't think I did it correctly
2. then, I got: X-3= (4 Square root of X+2)

Answer by checkley77(12844)

(Square root of 2X+3) = (square root of X+2) + 2
X-3=(4 Square root of X+2) GOOD SO FAR.
NOW SQUARE BOTH SIDES AGAIN.
X^2-6X+9=16(X+2)
X^2-6X+9=16X+32
X^2-6X-16X+9-32=0
X^2-22X-23=0
(X-23)(X+1)=0
X-23=0
X=23 ANSWER.
X+1=0
X=-1 ANSWER.
PROOF:
SQRT(2*23+3)=SQRT(23+2)+2
SQRT(46+3)=SQRT(25)+2
SQRT49=5+2
7=7